OpenStudy (anonymous):

I need to be able to explain in writing how I got the answer to square 64 divided by square 8...I know how to write it out but not explain it in writing how I got there

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{64^2}{8^2}=\frac{(8^2)^2}{8^2}=\frac{8^28^2}{8^2}=8^2=64\]

OpenStudy (anonymous):

there's pretty much nothing to write. tell 8^2=64 and that 64^2=64*64 and so wen they're divided u cancel 64 , resulting in 64.

OpenStudy (anonymous):

or: \[\frac{64^2}{8^2}=\left( \frac{64}{8} \right)^2=8^2=64\]

OpenStudy (anonymous):

or as brackett said